how can you rewrite this equation so that y is a function of x:
5(y-3x)=8-2x

Answer by nerdybill(5404)

5(y-3x)=8-2x
distributing the 5 to terms inside the parenthesis on the left:
5y-15x=8-2x
Adding 15x to both sides:
5y=13x+8
finally, dividing both sides by 5:
y = (13/5)x + (8/5)